If the hydrogen ion concentration of an unknown solution is 4.56E-7, what is its pH and will this form an acid or a base?


1
Expert's answer
2022-05-10T16:49:04-0400

Solution:

hydrogen ion = H+

[H+] = 4.56×10−7


pH = −log[H+]

pH = −log(4.56×10−7) = 6.34

pH = 6.34

pH < 7 acidic


Answer: pH = 6.34 (an acid)